OpenStudy (anonymous):

x^2-16 over x^2-2x-8

OpenStudy (anonymous):

factorize x^2-2x-8

OpenStudy (anonymous):

x = -2.4239 and x = 4.1510

OpenStudy (anonymous):

and also x^2-16

OpenStudy (shamil98):

\[\frac{ x^2-16 }{ x^2-2x-8 }\] \[= \frac{ (x-4)(x+4) }{ (x+4)(x-2) }\] Cancel the like terms to simplify.

OpenStudy (anonymous):

so it would be (x-4)(x-2) ?